# The life cycle of giant molecular clouds in simulated Milky Way-mass galaxies

### v1

@@ 2.4 Calculating GMC lifetimes and star formation efficiency
The lifetime of GMCs, which contains information on how fast molecular gas is consumed in galaxies, is crucial for understanding galactic star formation efficiency. While individual cloud lifetimes could, in principle, be estimated by tracking each cloud from its formation to dissolution, due to the complex mass assembly history of clouds in different environments, we use a Monte Carlo (MC) sampling method with probability-weighted tree paths (as described in Section 2.3) to give a simple characterization of the cloud population across different simulation runs.
In each MC iteration, we set one random walker per root node and initialize the lifetime as 00 Myr when the walker starts from the root node. Once a walker proceeds to the next node, i.e. it passes through one edge between two nodes, the lifetime of the current cloud tree increases by $1$ Myr, which corresponds to the temporal resolution of the constructed cloud network. In cases where a node has multiple descendants, the walker randomly chooses one descendant to proceed based on path probabilities calculated in Section 2.3. To fully explore the tree hierarchy, we perform $1000$ MC iterations to reach convergence for all model variations.
To associate the star formation activities (star particles) with the cloud life cycles (cloud merger trees), we propose the following approach. First, we define a ”main branch” for each root node, which refers to the path with the highest probability among all trajectories originating from that root node. Then star particles are matched to the main branches. For each node in any main branch, we record all gaseous cells’ IDs from the cloud of this node and check whether any of these IDs appear in star particles in the next simulation snapshot. If a star particle in the next snapshot shares an ID with one of the gas cells in this node, we identify that star particle as having formed from this node. We can associate most star particles and cloud merger trees well via such an approach. However, a corner case arises when a gas cell rapidly moves from cloud $A$ to cloud $B$ within the snapshot interval $\sim 1\mathrm{Myr}$, becomes a star particle within $B$, and is incorrectly attributed to cloud $A$. To address this issue, we add a distance limit: The matched star particle must be located within $10$ times the effective radius ($10\times R_{\mathrm{eff}}$) of its host cloud. We confirm that such cases are rare and that all results in this study remain unchanged if the distance limit is set to $5\times R_{\mathrm{eff}}$ instead of $10\times R_{\mathrm{eff}}$ as the distance limit.
After star particles and main branches are matched, we can calculate the integrated star formation efficiency of each main branch as
@@
where $t_{\mathrm{ff}}=\sqrt{3\pi/32G\bar{\rho}}$ is the gravitational free-fall timescale and the $\bar{\rho}$ is the average density over the volume of the cloud. Given that we determine the evolution tree at the finite time interval, the $\dot{M_{\star}}(t)$ is derived as
@@ 3 Results
Figure 1: Projections on the $x-y$ plane (face-on) for the gas surface density across the life cycle of one example GMC. The upper panel with three sub-panels shows respectively physical scales of $50$, $5$, and $0.5$ kpc on a side and illustrate the GMC’s position within the simulated disk galaxy at its birth. The bottom $9$ panels follow the evolution of this GMC at different ages as shown on the upper left of each panel. All star particles that form in this GMC are denoted with red stars. The black circle centered at the center-of-mass of the cloud on each panel indicates is the effective radius of the cloud at that time.
@@ 3.1 Visual impression of one GMC evolution
In total, we have constructed $3.9\times 10^{4}$, $9.0\times 10^{3}$, $9.1\times 10^{1}$, $8.3\times 10^{4}$, $3.1\times 10^{4}$, $1.4\times 10^{4}$ cloud merger trees respectively in the SFE1, SFE10, SFE100, Rad, SN, and Nofeed runs. In this section, we show one representative tree in the fiducial run as a case study to demonstrate how our algorithm successfully tracks the evolution of one single GMC in the simulated galaxies. In the upper panel of Fig. 1, the structure of one GMC identified in the fiducial run has been zoomed in from the large-scale galactic environment. The lower panel shows the evolutionary track of its life cycle. In the very early stages when we first identify the structure as the root node of the evolution tree (1-2 Myr), the size ($\sim 3\mathrm{pc}$), total gas mass ($\sim 10^{4}M_{\odot}$) and gas surface density ($\sim 10^{3}M_{\odot}\cdot\mathrm{pc}^{-2}$) of the cloud are small. During 2–4 Myr, as the cloud accretes gas from its surroundings, its mass and radius increase continuously. At $\sim 4\,\mathrm{Myr}$, the cloud has gaseous mass $\sim 2\times 10^{5}M_{\odot}$, a value close to the maximum mass reached during its life cycle, and virial parameter $\alpha_{\mathrm{vir}}\sim 0.6$, therefore starts to form stars. At $\sim 6\,\mathrm{Myr}$, the stellar mass within the GMC reaches its maximum $\sim 2\times 10^{4}M_{\odot}$. As young stars disturb the cloud through early feedback mechanisms, the virial parameter $\alpha_{\mathrm{vir}}$ increases slightly over unity. Finally, a considerable number of stars disperse the cloud at $\sim 8-9\,\mathrm{Myr}$ until we lose track of the cloud. Quantitatively, the maximum baryonic mass over the life cycle of the cloud is $\sim 2.2\times 10^{5}M_{\odot}$ and the integrated SFE $\varepsilon_{\mathrm{int}}$ is approximately $9$ percent according to equation (3).
Figure 2: Cumulative distribution of GMC mass ($M_{\mathrm{GMC}}$) in different model variations: SFE1 (royal blue), SFE10 (sky blue), SFE100 (light blue), Rad (red), SN (dark orange), and Nofeed (gray). Each solid line gives the median value of the cumulative mass function and the shaded areas enclose the $25\%-75\%$ percentile across the time span $0.4-1$ Gyr. We also show the slopes of $1.5$ and $2$ as black dashed lines for reference.
@@ 3.2 Physical properties of GMCs in different simulation runs
In Fig. 2, the GMC mass distribution above $10^{5}M_{\odot}$ in most runs except SFE100 and Nofeed has a power-law shape with a slope between -2 and -1.5, consistent with observations of Galactic GMCs (e.g. Fukui & Kawamura 2010). The mass function of the GMCs in the SFE100 run shows a steeper slope, indicating the depletion of massive GMCs with a mass greater than $10^{5}\,{\rm M_{\odot}}$. The extremely high $\epsilon_{\rm ff}$ in SFE100 leads to much faster consumption of cold gas and prevents clumps from growing into massive ones. It also leads to earlier and stronger stellar feedback that disperses the GMCs faster. On the other extreme, in the Nofeed run, due to the lack of stellar feedback, there exists a larger number of massive GMCs with a shallower slope compared to the fiducial run.
@@
Figure 4: 2D histograms of the velocity dispersion ($\sigma_{v}$) and the effective radius ($R_{\mathrm{eff}}$) of the identified clouds in all model variations labeled at the upper-left corner in each panel. We fit a Larson-like relation for each panel and the best-fit relation is over-plotted as black dashed lines. In the Nofeed run, the blue dashed line shows an additional fit to the data limited to the $\sigma_{v}<1\,\mathrm{km}\,\mathrm{s}^{-1}$ region.
@@ 3.3 GMC lifetime distribution
We then focus on the effects on cloud lifetimes from different channels of stellar feedback by comparing cloud lifetimes in the Rad run and the SN run with the fiducial run. In the Rad run, the cloud lifetime distribution follows almost the same slope as in the fiducial run for short-lived clouds with $t_{\mathrm{life}}<20\,\mathrm{Myr}$, suggesting that their evolution is mostly controlled by early radiative feedback rather than SN feedback. However, for $t_{\mathrm{life}}>20\,\mathrm{Myr}$, Rad shows an extended tail with a shallower slope compared with the fiducial run. This difference directly shows that many long-lived clouds cannot be disrupted by radiative feedback alone but have to wait for SN to quench the star formation activities. As we will discuss in Section 3.4 (see also Fig.  7), these long-lived clouds are systematically more massive and have higher gas surface density, than the short-lived ones. Therefore, the relatively mild photoionization and radiative pressure are too weak to terminate the lives of such massive clouds in the lack of sudden momentum and energy feedback from the supernova. On the opposite, we report the relatively shallower slope at the low lifetime end in the SN run. For these short-lived clouds, a supernova needs more time than the cloud’s lifetime to take effect so there is nearly no feedback during the life cycles of those short-lived clouds.
@@
To further address the physical nature of the two populations of clouds, we then separate the clouds in the fiducial run into two groups based on the distance of the clouds to the galactic center and plot the temporal evolution of $\alpha$ and $M$ in each group respectively. In Fig. 6, we find a distinct difference in the evolution of these two groups. The clouds beyond $2$ kpc from the galactic center go through life cycles as follows: The gaseous mass of these clouds is relatively low compared to the mass of clouds in the inner galactic region, e.g. $\sim 10^{4}$$10^{5}\,M_{\odot}$, and they are still accreting gas from the galactic environment. The virial parameters of these clouds generally start from a low value indicating the gravitationally bound state and then decrease during the accretion process. The gaseous masses correspondingly grow in this accreting stage. When it comes to the $1-2\mathrm{Myr}$ before the termination of clouds’ life cycles, the clouds are massive enough to collapse. After collapsing to form stars, clouds tend to be more gravitationally unbound due to the feedback mechanism and the virial parameters reach a much larger value. For the group of clouds closer to the galactic center, we however, see the quite different and significantly more disturbed trend of the virial parameters as the function of time. Compared with clouds in the inner region, these clouds are generally more massive and originate from unbound clumps. During their life cycles, they may remain marginally unbound due to the strong galactic scale effects on cloud dynamics by a combination of galactic shear and external pressure. The spatial separation is chosen as $2$ kpc here but we have checked that the above two modes exist similarly if the spatial separation is chosen as $1$, $1.5$, or $2.5$ kpc. We have also verified that the gas dynamics of clouds in the Nofeed, Rad, and SN runs all exhibit a similar spatial distribution trend.
@@ 3.4 Relation between GMC lifetime and maximum gaseous mass
In the bottom-left panel of Fig. 6, we observe that for clouds located outside the galaxy’s central region, the accretion process continuously supplements their gas content throughout most of their life cycle. Then they get disrupted quickly in a couple of million years and eventually, we lose track of them. Inspired by such a trend, we then analyze how the maximum gas mass of a cloud relates to its lifetime in Fig.  7. The discrete distribution in the lifetimes is due to the finite temporal resolution of our cloud evolution tree of $1$ Myr. In the SFE100 run, owing to the temporal resolution of the evolutionary tree, we could barely resolve the life cycles of the low-mass clouds due to the rapid gas consumption via the star formation process in these clouds. Despite the model variations in different simulation runs, all the lines show a significant positive correlation between the lifetimes and the gaseous masses in general, in agreement with the continuous accreting state of the GMCs in their early life. The lines for the Rad and SN runs both rise faster and are slightly higher than the line for the SFE1 run, while the Nofeed run corresponds to the line with the steepest slope across $\sim 4\times 10^{4}M_{\odot}$ to $\sim 5\times 10^{6}M_{\odot}$. Such a trend is consistent with the fact that both early feedback and SN feedback contribute to dispersing clouds. In most runs except SFE100, we find a rapid increase in the lifetime of GMCs in the low mass regime ($<10^{5}M_{\odot}$). But then the slope between the lifetime and gaseous mass becomes shallower in the $M_{\mathrm{max}}>10^{5}M_{\odot}$ regime. On the one hand, GMCs in their late lives may accrete gas from surrounding ISM not as efficiently as in their early lives. On the other hand, the star formation process continuously consumes the gas of GMCs. The combined effect of the above causes the different slopes between the lifetime and gaseous mass in different regimes.
@@ 3.5 Star formation efficiency on cloud scales
Recently, it has been realized from many theoretical and numerical works (e.g. Murray et al. 2010; Grudić et al. 2018; Li et al. 2019) that there exists a positive correlation between integrated star formation efficiency and gas surface density of clouds:
@@
where critical feedback density $\Sigma_{\mathrm{fb}}=\frac{\dot{p_{\star}}}{4\pi G}$ and $\dot{p_{\star}}$ is the time- and IMF-averaged rate of the specific momentum feedback strength from massive stars. The physical origin of this correlation is the force balance between the self-gravity of the cloud and momentum stellar feedback from massive stars (e.g. Fall et al. 2010; Li et al. 2019). This physical interpretation assumes that the cloud is in isolation without accreting material and without external forces from the ambient environment. Here we test this scenario using our galaxy simulations with resolved GMC populations.
Based on the method described in Section 2.4, we calculate the integrated SFE for each GMC along the main branch of the network. As we can see in Fig. 8, in general, GMCs in simulated galaxies follow a similar positive correlation between $\Sigma_{\rm max}$ and $\varepsilon_{\rm int}$ for all model variations. In the fiducial run, the correlation between $\Sigma_{\rm max}$ and $\varepsilon_{\rm int}$ is qualitatively in agreement with the isolated cloud studies (e.g. Grudić et al. 2018; Li et al. 2019) shown by equation (7). Comparing the fiducial and the SFE10 and SFE100 runs, we find that increasing $\epsilon_{\mathrm{ff}}$ leads to faster depletion of high gas surface density clouds because of high star formation on the cell level and thus earlier and stronger feedback. For lower $\Sigma_{\mathrm{max}}$ clouds, higher $\epsilon_{\mathrm{ff}}$ leads to higher $\varepsilon_{\rm int}$ meaning that the cloud-scale integrated efficiency follows the same trend of the cell-level efficiency. This is consistent with the trend found in the cosmological simulations (Agertz et al. 2013; Li et al. 2018; Oh et al. 2020; Nuñez-Castiñeyra et al. 2021). It is worth emphasizing here that the SFE100 is the most extreme case in which only a sparse number of clouds are identified, making them hard to track.
While the general distribution of $\varepsilon_{\mathrm{int}}$ is roughly consistent with equation (7) in the fiducial run, we would like to investigate how this relation varies when different channels of stellar feedback are used by comparing the GMC sample in the Rad, SN, and Nofeed runs. Unlike the fiducial run where stellar wind and radiation pressure from newly born stars prevent GMCs from collapsing to extremely high density in the early stage of star formation, in the SN run a considerable number of GMCs manage to grow into extremely dense clumps with $\Sigma_{\mathrm{max}}$ even close to $10^{4}\,M_{\odot}\mathrm{pc}^{-2}$ due to the lack of early stellar feedback. Clouds that reach high surface densities and collapse to form new stars can remain bound and continue star formation in the absence of early feedback until eventually being dispersed by supernovae from the older stars.
@@
In the Nofeed run, as expected, a considerable number of clouds with $\Sigma_{\mathrm{max}}>10^{3}\;M_{\odot}\mathrm{pc}^{-2}$ show extremely efficient star formation with the majority of the gaseous mass turning into stellar masses, i.e. reaching almost $100$ percent integrated star formation efficiency. Intriguingly, without any stellar feedback input, there are still a large number of clouds showing relatively low integrated star formation efficiency during their lifetime, especially for clouds with $\Sigma_{\mathrm{max}}<10^{3}\;M_{\odot}\mathrm{pc}^{-2}$, and the positive correlation between $\Sigma_{\mathrm{max}}$ and $\varepsilon_{\mathrm{int}}$ still roughly holds. This suggests that other mechanisms, possibly galactic dynamics, act as an effective “feedback source” that is responsible for the disruption of the GMCs.
Following this logic, we improve the model in equation (7) by adding a galactic shear term that takes into account the environmental effects over galactic scales. For simplicity, we introduce the critical shear surface density $\Sigma_{\mathrm{shear}}$ which effectively reflects how the galactic shear/tides suppress gravitational collapse. The equation (7) then becomes
@@
In the Nofeed run, as $\Sigma_{\mathrm{fb}}=0$ we can independently fit $\Sigma_{\mathrm{shear}}$ within our cloud sample. The best-fit value of the critical shear density is $\Sigma_{\mathrm{shear}}=3.9\times 10^{3}M_{\odot}\mathrm{pc}^{-2}$, which can be interpreted as a galaxy-wide average strength of the galactic shear/tides. Using this constant and fitting equation (8) to the SFE1 samples, we obtain the critical feedback density as $\Sigma_{\mathrm{fb}}=2.8\times 10^{3}M_{\odot}\mathrm{pc}^{-2}$. Although $\Sigma_{\mathrm{shear}}$ is not a first-principle measurement of galactic environmental effects on GMCs, the main takeaway from the above fit is that external galactic environments play a similar, if not bigger, role than internal stellar feedback in shaping GMC life cycles.

### v2

@@ 2.4 Calculating GMC lifetimes and star formation efficiency
The lifetime of GMCs, which contains information on how fast molecular gas is consumed in galaxies, is crucial for understanding galactic star formation efficiency. While individual cloud lifetimes could, in principle, be estimated by tracking each cloud from its formation to dissolution, due to the complex mass assembly history of clouds in different environments, we used a Monte Carlo (MC) sampling method with probability-weighted tree paths (as described in Section 2.3) to give a simple characterization of the cloud population across different simulation runs.
In each MC iteration, we set one random walker per root node and initialized the lifetime as 00 Myr when the walker starts from the root node. Once a walker proceeded to the next node, i.e., it passed through one edge between two nodes, the lifetime of the current cloud tree increased by $1$ Myr, which corresponded to the temporal resolution of the constructed cloud network. In cases where a node had multiple descendants, the walker randomly chose one descendant to proceed based on path probabilities calculated in Section 2.3. To fully explore the tree hierarchy, we performed $1000$ MC iterations to reach convergence for all model variations.
To associate the star formation activities (star particles) with the cloud life cycles (cloud merger trees), we proposed the following approach. First, we defined a ”main branch” for each root node, which referred to the path with the highest probability among all trajectories originating from that root node. Then star particles were matched to the main branches. For each node in any main branch, we recorded all gaseous cells’ IDs from the cloud of this node and checked whether any of these IDs appeared in star particles in the next simulation snapshot. If a star particle in the next snapshot shared an ID with one of the gas cells in this node, we identified that star particle as having formed from this node. We could associate most star particles and cloud merger trees well via such an approach. However, a corner case arised when a gas cell rapidly moved from cloud $A$ to cloud $B$ within the snapshot interval $\sim 1\,\mathrm{Myr}$, became a star particle within $B$, and was incorrectly attributed to cloud $A$. To address this issue, we added a distance limit: The matched star particle must be located within $10$ times the effective radius ($10\times R_{\mathrm{eff}}$) of its host cloud. We confirmed that such cases were rare and that all results in this study remained unchanged if the distance limit was set to $5\times R_{\mathrm{eff}}$ instead of $10\times R_{\mathrm{eff}}$ as the distance limit.
After star particles and main branches were matched, we calculated the integrated star formation efficiency of each main branch as
@@
where $t_{\mathrm{ff}}=\sqrt{3\pi/32G\bar{\rho}}$ is the gravitational free-fall timescale and the $\bar{\rho}$ is the average density over the volume of the cloud. Given that we determined the evolution tree at the finite time interval, the $\dot{M_{\star}}(t)$ was derived as
@@ 3 Results
Figure 1: Projections on the $x-y$ plane (face-on) for the gas surface density across the life cycle of one example GMC. The upper panel with three sub-panels shows respectively physical scales of $50$, $5$, and $0.5$ kpc on a side and illustrate the GMC’s position within the simulated disk galaxy at its birth. The bottom nine panels follow the evolution of this GMC at different ages as shown on the upper left of each panel. All star particles that form in this GMC are denoted with red stars. The black circle centered at the center-of-mass of the cloud on each panel indicates is the effective radius of the cloud at that time.
@@ 3.1 Visual impression of one GMC evolution
In total, we constructed $3.9\times 10^{4}$, $9.0\times 10^{3}$, $9.1\times 10^{1}$, $8.3\times 10^{4}$, $3.1\times 10^{4}$, $1.4\times 10^{4}$ cloud merger trees respectively in the SFE1, SFE10, SFE100, Rad, SN, and Nofeed runs. In this section, we show one representative tree in the fiducial run as a case study to demonstrate how our algorithm successfully tracks the evolution of one single GMC in the simulated galaxies. In the upper panel of Fig. 1, the structure of one GMC identified in the fiducial run has been zoomed in from the large-scale galactic environment. The lower panel shows the evolutionary track of its life cycle. In the very early stages when we first identify the structure as the root node of the evolution tree ($1-2$ $\mathrm{Myr}$), the size ($\sim 3$ $\mathrm{pc}$), total gas mass ($\sim 10^{4}$ $M_{\odot}$) and gas surface density ($\sim 10^{3}$ $M_{\odot}\cdot\mathrm{pc}^{-2}$) of the cloud are small. During $2-4$ $\mathrm{Myr}$, as the cloud accretes gas from its surroundings, its mass and radius increase continuously. At $\sim 4$ $\mathrm{Myr}$, the cloud has gaseous mass $\sim 2\times 10^{5}$ $M_{\odot}$, a value close to the maximum mass reached during its life cycle, and virial parameter $\alpha\sim 0.6$, therefore starts to form stars. At $\sim 6$ $\mathrm{Myr}$, the stellar mass within the GMC reaches its maximum $\sim 2\times 10^{4}$ $M_{\odot}$. As young stars disturb the cloud through early feedback mechanisms, the virial parameter $\alpha$ increases slightly over unity. Finally, a considerable number of stars disperses the cloud at $\sim 8-9$ $\mathrm{Myr}$ until we lose track of the cloud. Quantitatively, the maximum baryonic mass over the life cycle of the cloud is $\sim 2.2\times 10^{5}$ $M_{\odot}$ and the integrated SFE $\varepsilon_{\mathrm{int}}$ is approximately $9\%$ according to equation (7).
Figure 2: Cumulative distribution of GMC mass ($M_{\mathrm{GMC}}$) in different model variations: SFE1 (royal blue), SFE10 (sky blue), SFE100 (light blue), Rad (red), SN (dark orange), and Nofeed (gray). Each solid line gives the median value of the cumulative mass function and the shaded areas enclose the $25\%-75\%$ percentile across the time span $0.4-1$ Gyr. We also show the slopes of $1.5$ and $2$ as dashed black lines for reference.
@@ 3.2 Physical properties of GMCs in different simulation runs
In Fig. 2, the GMC mass distribution above $10^{5}M_{\odot}$ in most runs except SFE100 and Nofeed has a power-law shape with a slope between -2 and -1.5, consistent with observations of Galactic GMCs (e.g., Fukui & Kawamura 2010). The mass function of the GMCs in the SFE100 run shows a steeper slope, indicating the depletion of massive GMCs with a mass greater than $10^{5}\,{\rm M_{\odot}}$. The extremely high $\epsilon_{\rm ff}$ in SFE100 leads to much faster consumption of cold gas and prevents clumps from growing into massive ones. It also leads to earlier and stronger stellar feedback that disperses the GMCs faster. Consequently, less massive GMCs in SFE100 tend to be weakly bound and have larger $\alpha$ (see Fig. 3). On the other extreme, in the Nofeed run, due to the lack of stellar feedback, there exists a larger number of massive GMCs with a shallower slope compared to the fiducial run.
@@
Figure 4: 2D histograms of the velocity dispersion ($\sigma_{v}$) and the effective radius ($R_{\mathrm{eff}}$) of the identified clouds in all model variations labeled at the upper left corner in each panel. We fit a Larson-like relation for each panel and the best-fit relation is over-plotted as dashed black lines. In the Nofeed run, the dashed blue line shows an additional fit to the data limited to the $\sigma_{v}<1\,\mathrm{km}\,\mathrm{s}^{-1}$ region.
@@ 3.3 GMC lifetime distribution
We then focus on the effects on cloud lifetimes from different channels of stellar feedback by comparing cloud lifetimes in the Rad run and the SN run with the fiducial run. In the Rad run, the cloud lifetime distribution follows almost the same slope as in the fiducial run for short-lived clouds with $t_{\mathrm{life}}<20\,\mathrm{Myr}$, suggesting that their evolution is mostly controlled by early feedback rather than SN feedback. However, for $t_{\mathrm{life}}>20\,\mathrm{Myr}$, Rad shows an extended tail with a shallower slope compared with the fiducial run. This difference directly shows that many long-lived clouds cannot be disrupted by early feedback alone but have to wait for SN to quench the star formation activities. As we discuss in Section 3.4 (see also Fig.  7), these long-lived clouds are systematically more massive and have higher gas surface density, than the short-lived ones. Therefore, the relatively mild photoionization and radiative pressure are too weak to terminate the lives of such massive clouds in the lack of sudden momentum and energy feedback from the supernova. On the opposite, we report the shallower slope at the low lifetime end in the SN run. For these short-lived clouds, a supernova needs more time than the cloud’s lifetime to take effect so there is nearly no feedback during the life cycles of those short-lived clouds.
@@
To further address the physical nature of the two populations of clouds, we then separate the clouds in the fiducial run into two groups based on the distance of the clouds to the galactic center and plot the temporal evolution of $\alpha$ and $M$ in each group respectively. In Fig. 6, we find a distinct difference in the evolution of these two groups. The clouds beyond $2$ kpc from the galactic center go through life cycles as follows: The gaseous mass of these clouds is relatively low compared to the mass of clouds in the inner galactic region, e.g., $\sim 10^{4}$$10^{5}\,M_{\odot}$, and they are still accreting gas from the galactic environment. The virial parameters of these clouds generally start from a low value indicating the gravitationally bound state and then decrease during the accretion process. The gaseous masses correspondingly grow in this accreting stage. When it comes to the $1-2\,\mathrm{Myr}$ before the termination of clouds’ life cycles, the clouds are massive enough to collapse. After collapsing to form stars, clouds tend to be more gravitationally unbound due to the feedback mechanism and the virial parameters reach a much larger value. For the group of clouds closer to the galactic center, we however, see the quite different and significantly more disturbed trend of the virial parameters as the function of time. Compared with clouds in the inner region, these clouds are generally more massive and originate from unbound clumps. During their life cycles, they may remain marginally unbound due to the strong galactic scale effects on cloud dynamics by a combination of galactic shear and external pressure. The spatial separation is chosen as $2$ kpc here but we have checked that the above two modes exist similarly if the spatial separation is chosen as $1$, $1.5$, or $2.5$ kpc. We have also verified that the gas dynamics of clouds in the Nofeed, Rad, and SN runs all exhibit a similar spatial distribution trend.
@@ 3.4 Relation between GMC lifetime and maximum gaseous mass
In the bottom left panel of Fig. 6, we observe that for clouds located outside the galaxy’s central region, the accretion process continuously supplements their gas content throughout most of their life cycle. Then they get disrupted quickly in a couple of million years and eventually, we lose track of them. Inspired by such a trend, we then analyze how the maximum gas mass of a cloud relates to its lifetime in Fig.  7. The discrete distribution in the lifetimes is due to the finite temporal resolution of our cloud evolution tree of $1$ Myr. In the SFE100 run, owing to the temporal resolution of the evolutionary tree, we could barely resolve the life cycles of the low-mass clouds due to the rapid gas consumption via the star formation process in these clouds. Despite the model variations in different simulation runs, all the lines show a significant positive correlation between the lifetimes and the gaseous masses in general, in agreement with the continuous accreting state of the GMCs in their early life. The lines for the Rad and SN runs both rise faster and are slightly higher than the line for the SFE1 run, while the Nofeed run corresponds to the line with the steepest slope across $\sim 4\times 10^{4}M_{\odot}$ to $\sim 5\times 10^{6}M_{\odot}$. Such a trend is consistent with the fact that both early feedback and SN feedback contribute to dispersing clouds. In most runs except SFE100, we find a rapid increase in the lifetime of GMCs in the low mass regime ($<10^{5}M_{\odot}$). But then the slope between the lifetime and gaseous mass becomes shallower in the $M_{\mathrm{max}}>10^{5}M_{\odot}$ regime. On the one hand, GMCs in their late lives may accrete gas from surrounding ISM not as efficiently as in their early lives. On the other hand, the star formation process continuously consumes the gas of GMCs. The combined effect of the above causes the different slopes between the lifetime and gaseous mass in different regimes.
@@ 3.5 Star formation efficiency on cloud scales
Recently, it has been realized from many theoretical and numerical works (e.g., Murray et al. 2010; Grudić et al. 2018; Li et al. 2019) that there exists a positive correlation between integrated star formation efficiency and gas surface density of clouds:
@@
where critical feedback density $\Sigma_{\mathrm{fb}}=\frac{\dot{p_{\star}}}{4\pi G}$ and $\dot{p_{\star}}$ is the time- and IMF-averaged rate of the specific momentum feedback strength from massive stars. The physical origin of this correlation is the force balance between the self-gravity of the cloud and momentum stellar feedback from massive stars (e.g., Fall et al. 2010; Li et al. 2019). This physical interpretation assumes that the cloud is in isolation without accreting material and without external forces from the ambient environment. Here we test this scenario using our galaxy simulations with resolved GMC populations.
Based on the method described in Section 2.4, we calculate the integrated SFE for each GMC along the main branch of the network. As we can see in Fig. 8, in general, GMCs in simulated galaxies follow a similar positive correlation between $\Sigma_{\rm max}$ and $\varepsilon_{\rm int}$ for all model variations. In the fiducial run, the correlation between $\Sigma_{\rm max}$ and $\varepsilon_{\rm int}$ is qualitatively in agreement with the isolated cloud studies (e.g., Grudić et al. 2018; Li et al. 2019) shown by equation (11). Comparing the fiducial and the SFE10 and SFE100 runs, we find that increasing $\epsilon_{\mathrm{ff}}$ leads to faster depletion of high gas surface density clouds because of high star formation on the cell level and thus earlier and stronger feedback. For lower $\Sigma_{\mathrm{max}}$ clouds, higher $\epsilon_{\mathrm{ff}}$ leads to higher $\varepsilon_{\rm int}$ meaning that the cloud-scale integrated efficiency follows the same trend of the cell-level efficiency. This is consistent with the trend found in the cosmological simulations (Agertz et al. 2013; Li et al. 2018; Oh et al. 2020; Nuñez-Castiñeyra et al. 2021). It is worth emphasizing here that the SFE100 is the most extreme case in which only a sparse number of clouds are identified, making them hard to track.
While the general distribution of $\varepsilon_{\mathrm{int}}$ is roughly consistent with equation (11) in the fiducial run, we would like to investigate how this relation varies when different channels of stellar feedback are used by comparing the GMC sample in the Rad, SN, and Nofeed runs. Unlike the fiducial run where stellar wind and radiation pressure from newly born stars prevent GMCs from collapsing to an extremely high density in the early stage of star formation, in the SN run a considerable number of GMCs manage to grow into extremely dense clumps with $\Sigma_{\mathrm{max}}$ even close to $10^{4}\,M_{\odot}\mathrm{pc}^{-2}$ due to the lack of early stellar feedback. Clouds that reach high surface densities and collapse to form new stars can remain bound and continue star formation in the absence of early feedback until eventually being dispersed by supernovae from the older stars.
@@
In the Nofeed run, as expected, a considerable number of clouds with $\Sigma_{\mathrm{max}}>10^{3}\;M_{\odot}\mathrm{pc}^{-2}$ show extremely efficient star formation with the majority of the gaseous mass turning into stellar masses, i.e., reaching almost $100\%$ integrated star formation efficiency. Intriguingly, without any stellar feedback input, there are still a large number of clouds showing relatively low integrated star formation efficiency during their lifetime, especially for clouds with $\Sigma_{\mathrm{max}}<10^{3}\;M_{\odot}\mathrm{pc}^{-2}$, and the positive correlation between $\Sigma_{\mathrm{max}}$ and $\varepsilon_{\mathrm{int}}$ still roughly holds. This suggests that other mechanisms, possibly galactic dynamics, act as an effective “feedback source” that is responsible for the disruption of the GMCs.
Following this logic, we improve the model in equation (11) by adding a galactic shear term that takes into account the environmental effects on the galactic scale. For simplicity, we introduce the critical shear surface density $\Sigma_{\mathrm{shear}}$ which effectively reflects how the galactic shear/tides suppress gravitational collapse. The equation (11) then becomes
@@
In the Nofeed run, as $\Sigma_{\mathrm{fb}}=0$ we can independently fit $\Sigma_{\mathrm{shear}}$ within our cloud sample. The best-fit value of the critical shear density is $\Sigma_{\mathrm{shear}}=3.9\times 10^{3}M_{\odot}\mathrm{pc}^{-2}$, which can be interpreted as a galaxy-wide average strength of the galactic shear and tides. Using this constant and fitting equation (12) to the SFE1 samples, we obtain the critical feedback density as $\Sigma_{\mathrm{fb}}=2.8\times 10^{3}M_{\odot}\mathrm{pc}^{-2}$. Although $\Sigma_{\mathrm{shear}}$ is not a first-principle measurement of galactic environmental effects on GMCs, the main takeaway from the above fit is that external galactic environments play a similar, if not bigger, role than internal stellar feedback in shaping GMC life cycles.